Explain how to make up 600 ml of 0.2 M sodium hydroxide solution.
1
Expert's answer
2020-01-31T04:35:06-0500


Explain how to make up 600 ml of 0.2 M sodium hydroxide solution.


Solution:

1) Sodium hydroxide - NaOH

Molecular weight of sodium hydroxide = M(NaOH) = Ar(Na) + Ar(O) + Ar(H) = 23 + 16 + 1 = 40 (g/mol).

2) Molarity (M) is the concentration of a solution expressed as the number of moles of solute per liter of solution:

Molarity (M) = moles solute / liters solution = CM = n / V;

3) n = m / M, then:

CM = n / V = m /(M*V).

Then,

m (NaOH) = Cm * M * V = 0.2 M * 40 g/mol * 0.6 L = 4.8 g.

m(NaOH) = 4.8 grams.


Answer: Add 4.8 g NaOH to 600mL of deionized water, you will end up with 600mL of 0.2M NaOH solution.